### v1

@@ Abstract
We develop an information-theoretic framework for bounding the expected supremum and tail probabilities of stochastic processes, offering a simpler and sharper alternative to classical chaining and slicing arguments for generalization bounds. The key idea is a lifting argument that produces information-theoretic analogues of empirical process bounds, such as Dudley’s entropy integral. The lifting introduces symmetry, yielding sharp bounds even when the classical Dudley integral is loose. As a by-product, we obtain a concise proof of the majorizing measure theorem, providing explicit constants. The information-theoretic approach provides a soft version of classical localized complexity bounds in generalization theory, but is more concise and does not require the slicing argument. We apply this approach to empirical risk minimization over Sobolev ellipsoids and weak $\ell_{q}$ balls, obtaining sharper convergence rates or extensions to settings not covered by classical methods.
@@ 1 Introduction
Sharp control of empirical processes is central to modern statistical learning theory, underpinning generalization bounds, minimax rates, and risk analysis for high-dimensional estimators. Classical tools such as Dudley’s entropy integral have played a central role in this analysis, with applications ranging from nonparametric regression (Van de Geer,, 1990; Wainwright,, 2019) to sparse high-dimensional models (Raskutti et al.,, 2011). Although certain special cases can be handled by simpler techniques, the localized version of Dudley’s integral (Van de Geer,, 1990) has remained a fundamental tool in many general settings. However, its application to statistical learning may still be limited since (i) Dudley’s integral, while widely used, can be suboptimal even for basic examples such as Sobolev ellipsoids, and (ii) the slicing (or peeling) technique for localization often requires delicate covering number estimates, sometimes invoking deep results from approximation theory or the local theory of Banach spaces, e.g. in sparse regression over the $\ell_{q}$ ball (Raskutti et al.,, 2011).
In this paper, we introduce a “lifting” method for deriving new information-theoretic generalization bounds, which overcomes the limitations discussed above. The key idea is to introduce replicated copies of a system and apply the method of types from information theory, so that traditional “hard” inequalities for the supremum of a stochastic process for an index set yield “soft” information-theoretic inequalities for a distribution on that set. The strength of the lifting approach lies in leveraging geometric insights involving covering and packing numbers, such as the chaining argument in Dudley’s integral, yielding new information-theoretic inequalities that are not easily derived from traditional information-theoretic methods. This method yields sharper localized generalization bounds and new results for empirical risk minimization over ellipsoids and sparse regression.
@@
where $\Pi_{\sigma}(\mu)$ denotes the set of couplings of $\mu$ and itself with mean square error $\sigma^{2}$, I(Z;Z^)I($\hat{Z}$) is the mutual information, and $R_{\mu}(\sigma^{2})$ denotes the rate-distortion function (Thomas and Joy,, 2006). The supremum in (1) is over all couplings of the index and the process satisfying the marginal condition $P_{Z}=\mu$. While the analogy between the rate-distortion function $R_{\mu}(\sigma^{2})$ and covering numbers is well-known (Thomas and Joy,, 2006; Csiszár and Körner,, 2011), we are not aware of any prior literature explicitly stating (1). Previously, similar lifting arguments were used for problems in selected problems from network information theory (Wu et al.,, 2018; Liu,, 2022, 2021, 2023), but the application to empirical processes appears to be new. More generally, “soft” versions of traditional measure concentration techniques are known to provide sharper convergence in some hypothesis testing and information theory problems (Liu,, 2018; Liu et al.,, 2020; Liu,, 2020; Liu and Özgür,, 2020).
@@
First, our approach yields sharper convergence rate in some basic estimation problems. The reason is that (1) is a sharp two-sided inequality for any $\mu$, which might be surprising since we derive (1) from the Dudley integral, and the latter is not sharp for some $T$. The reason for the sharpness of (1) is that the lifting argument transforms the process into a stationary one - i.e., invariant under a transitive group of permutations - even if the original process lacks this symmetry. Since Dudley’s integral is sharp for stationary processes (see Section 2.1 for background), (1) inherits the sharpness, even if the original process is nonstationary. The sharpness of (1) also has an interesting consequence: by essentially maximizing over $\mu$ and applying the date processing inequality, we obtain a concise new proof of the nontrivial lower bound part of the classical majorizing measure theorem. This new proof of the majorizing measure theorem is conceptually simple, as it closely mirrors the argument used for the Dudley integral upper bound, essentially reversing its key steps. Furthermore, many nuisance parameters in the traditional proof disappear due to taking limits in the lifting argument. Figure 1 illustrates the implication relations of these results.
@@
Figure 1: Arrows indicate implications of results, and the techniques/assumptions for the implications. For Gaussian processes, the majorizing measure theorem and the proposed rate-distortion integral are two-sided bounds. Dudley’s integral is only an upper bound, unless stationarity is assumed.
@@
Second, our approach makes localization of generalization bounds more straightforward. Localization refers to the phenomenon where the index $Z$ of the process concentrates on a subset much smaller than the full index set $T$, so that taking the supremum over all of $T$ may significantly overestimate the typical value of the process at $Z$. For example, in empirical risk minimization (ERM), $Z$ denotes the estimator, which tends to concentrate near zero as the sample size increases. Traditionally, localization is achieved by the “slicing” (or “peeling”) argument, which yields sharp convergence rate for some empirical risk minimization problems Wainwright, (2019). Slicing typically involves controlling the modulus of continuity of the process through careful controls of the deviation magnitude for different slices. In contrast, using properties of the rate-distortion function, we can often directly upper bound the right side of (1) using the variance of $Z\sim\mu$, which is much smaller than the size of $T$. Similarly, tail probability bounds on $X_{Z}$, based on the moment generating function of $\|Z\|_{2}$, can also be derived using a lifting technique. This new method of localization does not require slicing.
To demonstrate our approach, we derive new and sharper statistical rates in two classical ERM problems:
  • For mean estimation using ERM over the Sobolev ellipsoid with smoothness parameter $\beta>1/2$, which plays a fundamental role in nonparametric statistics (Tsybakov,, 2009), the slicing method based on Dudley’s integral only gives an upper bound $O(\frac{1}{2\beta-1})n^{-\frac{2\beta}{2\beta+1}}$ for the mean square error. The factor $\frac{1}{2\beta-1}$ arises in Dudley’s integral for the supremum over the Sobolev ellipsoid, a classical example for the looseness of Dudley’s integral (Talagrand,, 2014). On the other hand, our method achieves the optimal $O(\frac{1}{\sqrt{2\beta-1}})n^{-\frac{2\beta}{2\beta+1}}$ bound.
  • For sparse regression over the $\ell_{q}$-ball, $q\in(0,1]$, Raskutti et al. (Raskutti et al.,, 2011) showed that the mean square prediction error is upper bounded by $(\frac{\ln d}{n})^{\frac{2-q}{2}}R$ up to a constant depending on $q$, where $d$ is the feature dimension, $n$ is the sample size, $R^{1/q}$ is the $\ell_{q}$ radius, the noise is standard normal, and the columns of the design matrix are normalized to have $\ell_{2}$ norm at most $\sqrt{n}$. The proof in Raskutti et al., (2011) was based on the slicing method and estimating covering numbers of the $q$-absolutely convex hull, using deep results from convex geometry (Guedon and Litvak,, 2000; Carl and Pajor,, 1988); see proof of (Raskutti et al.,, 2011, Lemma 3). We provide a new proof without citing the convex geometry literature, using only simple properties of the entropy, such as the chain rule, to extend elementary one-dimensional estimates to higher dimensions. In fact, we show that the same convergence rate holds for weak $\ell_{q}$ balls ($q\in(0,1)$), which is new and appears to be beyond the reach of prior techniques in the literature.
@@
Related work: A closely related line of recent research, started by Russo and Zou, (2016) and further developed by other authors including Xu and Raginsky, (2017); Dziugaite and Roy, (2017); Asadi et al., (2018); Bu et al., (2020); Aminian et al., (2023), proved information-theoretic generalization bounds in terms of the mutual information between the random index and the process. A basic version (see for example Russo and Zou, (2016)) is
$\mathbb{E}[X_{Z}]\lesssim\sup_{t\in T}\|X_{t}\|_{\Psi_{2}}\sqrt{I(Y;Z)}$ (2)
@@
where the sub-Gaussian norm $\|\cdot\|_{\Psi_{2}}$ will be defined in (6), and $Y:=(X_{t})_{t\in T}$ denotes the stochastic process (in the case of finite $T$, we have $Y\in\mathbb{R}^{|T|}$). While this can be interpreted as a strengthening of a maximal inequality (Jiao et al.,, 2018), further refinement using the chaining argument, strengthening the Dudley integral, is also possible (Asadi et al.,, 2018). This type of results are useful in statistics and machine learning, where $Z$ represents the output of an algorithm which has limited mutual information with the training data, potentially much smaller than the covering number of $T$. Though also concerning a random index framework, their results differ from and are not directly comparable with (1), since their strengthening is in terms of the mutual information between $Z$ and the data, rather than the rate-distortion function of $Z$. Their proofs typically involve information-theoretic arguments such as chain rules and nonnegativity of the relative entropy, but not the lifting technique in the present paper. Nevertheless, we will show in Theorem 2 a result that encompasses both (1) and (2), highlighting the versatility of the lifting technique.
In related work on empirical risk minimization (ERM), it has been shown that under favorable conditions, such as convex regression classes with Gaussian noise, the mean-squared error (MSE) can often be characterized by a fixed-point equation involving a localized Gaussian width (Chatterjee,, 2014). For ellipsoids, this localized Gaussian width can then be controlled directly via the Cauchy–Schwarz inequality (Wei et al.,, 2020), thereby avoiding the slicing and Dudley-integral machinery and mitigating the looseness of Dudley’s bound. However, this shortcut relies essentially on concentration of the error, which can fail when the regression class is nonconvex (e.g., sparse regression under the $\ell_{q}$ constraint with $q<1$) or when the noise is sub-Gaussian but does not satisfy a measure concentration inequality (such as log-Sobolev or transportation-cost inequality). Moreover, the Cauchy-Schwarz step exploits linear structure: for example, consider a regression class given by the image of a Sobolev ellipsoid under a Lipschitz map (e.g., induced by a nonlinear neural network). This generalizes the setting of Wei et al., (2020), but their argument no longer applies. These limitations are expected to be artifacts of the proof: tail probability bounds need not require concentration (as naturally follows from the Dudley integral; see Lemma 2), and Lipschitz maps do not increase the supremum of the relevant stochastic process (a consequence of the majorizing measure theorem; see (13)). Unfortunately, existing tools do not suffice to rigorously establish these intuitions, and our work introduces an alternative framework that overcomes these limitations.
Organization: The paper is organized as follows: Section 2 provides more background on upper bounds for stochastic processes from probability theory, the traditional slicing technique for ERM, and information-theoretic tools used in our proof. Section 3 presents our main results, including (1), its sharpness for Gaussian processes, a simple proof of the classical majorizing measure theorem based on (1), a general framework for error tail bounds in ERM including results on the ellipsoid and sparse regression examples mentioned above. The remainder consists of details and proofs supporting the main results.
@@ 2 Preliminaries
Notation: We use $\wedge$ and $\vee$ to denote the max and min of two numbers, and $x]_{+}:=x\vee 0$. Given probability measures $P$ and $Q$ on a metric space $(T,d)$, $W_{p}(P,Q):=\sup_{P_{XY}\in\Pi(P,Q)}\mathbb{E}^{\frac{1}{p}}[d(X,Y)^{p}]$ denotes the Wasserstein distance, where $\Pi(P,Q)$ denotes the set of couplings. For brevity, $\Pi(P):=\Pi(P,P)$. Product measures, such as $P\times Q$ or $P^{\otimes N}$, indicate independence. ${\rm diam}(T):=\sup_{x,y\in T}d(x,y)$ denotes the diameter of a set $T$. Standard Landau notations such as $o(\dots)$ are used. $\lesssim_{q}$ means less than or equal up to a constant depending on $q$, and $\asymp_{q}$ means $\lesssim_{q}$ and $\gtrsim_{q}$ are both true. For any $q>0$, the $\ell_{q}$ ball and the weak $\ell_{q}$ ball of radius $r$ centered at 0 are
$B_{q}(r)$ $:=\{x\in\mathbb{R}^{d}\colon\sum_{i=1}^{d}x_{i}^{q}\leq r^{q}\},$ (3)
$wB_{q}(r)$ $:=\{x\in\mathbb{R}^{d}\colon\sup_{t>0}t^{q}\sum_{i=1}^{d}1_{|x_{i}|>t}\leq r^{q}\}.$ (4)
More generally, $B_{q}(z,r):=z+B_{q}(r)$ if the center is $z$. For brevity, $B(z,r):=B_{2}(z,r)$. $\ell_{2}$ denotes the Hilbert space of square-summable infinite sequences. The empirical distribution of a sequence $x=(x(1),x(2),\dots,x(d))$ is denoted by $\widehat{P}_{x}$. Given a set $T$ in $\mathbb{R}^{d}$ and a matrix $A\in\mathbb{R}^{n\times d}$, define $T-T:=\{t-s\colon t,s\in T\}$ and $AT:=\{At\colon t\in T\}$.
@@ 2.1 Upper Bounds for Stochastic Processes
Classically, Dudley’s integral provides upper bounds on a sub-Gaussian process in terms of the covering numbers. Let $(T,d)$ be a metric space, and suppose that $(X_{t})_{t\in T}$ is a centered sub-Gaussian process, that is, $\mathbb{E}[X_{t}]=0$, and the natural metric
$d(t,t^{\prime}):=\|X_{t}-X_{t^{\prime}}\|_{\Psi_{2}},\quad\forall t,t^{\prime}\in T,$ (5)
where, following Vershynin, (2018), we defined the sub-Gaussian norm
$\|X\|_{\Psi_{2}}=\inf\{K>0\colon\mathbb{E}[e^{X^{2}/K^{2}}]\leq 2\}.$ (6)
Dudley’ integral states that
$\mathbb{E}[\sup_{t\in T}X_{t}]\leq K\int_{0}^{\Delta}\sqrt{\ln N(T,\lambda)}d\lambda$ (7)
@@
where $K$ is a universal constant, $N(T,\lambda)$ denotes the $\lambda$-covering number of $T$, i.e. the minimal number of (open) balls of radius $\lambda$ required to cover $T$, and $\Delta$ denotes the diameter of $T$. While usually called Dudley’s integral, (7) is occasionally attributed to other authors as well; see Dudley, (2016); Maurey, (2024); Talagrand, (2014) for more about the history. In the case of Gaussian processes, the natural metric is defined as
$d(t,t^{\prime}):=\sqrt{\mathbb{E}[|X_{t}-X_{t^{\prime}}|^{2}]}.$ (8)
Note that (8) differs from (5) by a universal constant factor (Vershynin,, 2018). Throughout this paper, when we discuss Gaussian processes specifically, we define the natural metric by (8) rather than (5). In the Gaussian case, a commonly referenced explicit bound is $K<24$ (Ledoux and Talagrand,, 2013). As pointed out in (Dudley,, 2016, Section 5.3), better estimates $K$ exist in the literature, and the proof essentially relies on the estimate
$\mathbb{E}[\max_{1\leq j\leq N}X_{j}]\leq\sqrt{2\ln N}$ (9)
where $X_{1},\dots,X_{N}$ are centered, jointly Gaussian, with variances not exceeding 1. If we drop Gaussianity and instead assume that $\|X_{1}\|_{\Psi_{2}},\dots,\|X_{N}\|_{\Psi_{2}}\leq 1$, then $\mathbb{E}[\max_{1\leq j\leq N}X_{j}]\leq\sqrt{2\ln(2N)}$ for $N\geq 1$ (Vershynin,, 2018, Section 2.7.3), which is slightly worse than the Gaussian case (9). We will see, however, that the difference will vanish when we apply the lifting argument, which effectively takes $N\to\infty$.
@@
For stationary Gaussian processes, it is known since Fernique, (1975) that Dudley’s integral is tight:
$\int_{0}^{\Delta}\sqrt{\ln N(T,\lambda)}d\lambda\leq K^{\prime}\mathbb{E}[\sup_{t\in T}X_{t}]$ (10)
@@
for a universal constant $K^{\prime}>0$. Maurey, (2024) gives a bound of $K^{\prime}\leq 432$. In the stationary case, the smaller balls in the chaining argument are equivalent in each level, so the chaining process in the proof of the Dudley integral can be reversed using the so-called super-Sudakov minoration (see Lemma 10 ahead).
@@
For general nonstationary processes, the same paper of Fernique (Fernique,, 1975) conjectured what was later known as the majorizing measure theorem, whose nontrivial lower bound part was proved more than a decade later by Talagrand (Talagrand,, 1987) using a more sophisticated generic chaining argument. Given a probability measure $\mu$ on $T$, define
$I_{\mu}(t):=\int_{0}^{\Delta}\sqrt{\ln\frac{1}{\mu(B(t,\lambda))}}d\lambda$ (11)
@@
The usual definition of $\gamma_{2}(T)$ in terms of admissible partitions (cf. Talagrand, (2014)) is equivalent to (12) up to a constant factor. Generally speaking, $\gamma_{2}(T)$ is more difficult to evaluate than Dudley’s integral: $\mu$ needs to be optimized for a quantity depending on $\mu(B(t,\lambda))$ over different scales of $\lambda$, whereas in Dudley’s integral, we can calculate the covering number by optimizing for each scale separately.
@@
The special case where $(X_{t})$ is also Gaussian is called the Gaussian comparison inequality, usually proved by an interpolation argument which is highly Gaussian-specific (Ledoux and Talagrand,, 2013). The more general sub-Gaussian case can be immediately seen from the majorizing measure theorem. As noted in Van Handel, (2014), this version of sub-Gaussian comparison is a nontrivial consequence of the majorizing measure theorem, for which no other proof is known.
It is also known (see Talagrand, (2014)) that the following quantities are equivalent to $\gamma_{2}(T)$ up to constants:

### v3

@@ Abstract
We develop an information-theoretic framework for bounding the supremum of stochastic processes, offering a simpler and sharper alternative to classical chaining and slicing arguments for generalization bounds. The key idea is a lifting argument that produces information-theoretic analogues of empirical process bounds, such as Dudley’s entropy integral. Lifting introduces permutation symmetry, yielding sharp bounds when the classical Dudley integral is loose. This gives a simple proof of the majorizing measure theorem via the sharpness of Dudley’s entropy integral for stationary processes, a result known well before the proof of the majorizing measure theorem. Furthermore, the information-theoretic formulation provides soft versions of classical localized complexity bounds in generalization theory, but is simpler and does not require the slicing argument. We apply this approach to empirical risk minimization over Sobolev ellipsoids, obtaining sharp convergence rates in settings where previous methods are suboptimal.
@@ 1 Introduction
Sharp control of empirical processes is central to modern statistical learning theory, underpinning generalization bounds, minimax rates, and risk analysis for high-dimensional estimators. Classical tools such as Dudley’s entropy integral have played a central role in this analysis, with applications ranging from nonparametric regression [41, 44] to sparse high-dimensional models [35]. Although certain special cases can be handled by simpler techniques, the localized version of Dudley’s integral [41] has remained a fundamental tool in many general settings. However, its application to statistical learning may still be limited since (i) Dudley’s integral, while widely used, can be suboptimal even for basic examples such as Sobolev ellipsoids, and (ii) the slicing (or peeling) technique for localization often requires delicate covering number estimates, sometimes invoking deep results from approximation theory or the local theory of Banach spaces, e.g. in sparse regression over the $\ell_{q}$ ball [35].
In this paper, we introduce a “lifting” method for deriving new information-theoretic generalization bounds, which overcomes the limitations discussed above. The key idea is to introduce replicated copies of a system and apply the method of types from information theory, so that traditional “hard” inequalities for the supremum of a stochastic process for an index set yield “soft” information-theoretic inequalities for a distribution on that set. The strength of the lifting approach lies in leveraging geometric insights involving covering and packing numbers, such as the chaining argument in Dudley’s integral, yielding new information-theoretic inequalities that are not easily derived from traditional information-theoretic methods. This method yields sharper localized generalization bounds and new results for empirical risk minimization in some settings.
@@
where $\Pi_{\sigma}(\mu)$ denotes the set of couplings of $\mu$ and itself with mean squared error $\sigma^{2}$, I(Z;Z^)I($\hat{Z}$) is the mutual information, and $R_{\mu}(\sigma^{2})$ denotes the rate-distortion function (Definition 3). The supremum in (1) is over all conditional distributions of the index $Z$ given the process such that $P_{Z}=\mu$. While the analogy between the rate-distortion function $R_{\mu}(\sigma^{2})$ and covering numbers is well-known [12, 13], we are not aware of any prior literature explicitly stating (1). Previously, similar lifting arguments were used for problems in selected problems from network information theory [46, 26, 25, 27], but the application to empirical processes appears to be new. More generally, “soft” versions of traditional measure concentration techniques are known to provide sharper convergence in some hypothesis testing and information theory problems [23, 30, 24, 28].
@@
First, our approach yields sharper convergence rate in some basic estimation problems. The reason is that (1) is a sharp two-sided inequality for any $\mu$, which is perhaps surprising, since (1) is derived from the Dudley integral, and the latter is not sharp for some $T$. The reason for the sharpness of (1) is that the lifting argument transforms a general Gaussian process into a stationary one (i.e., invariant under the transitive group of permutations; see Definition 1), even if the original process lacks this symmetry. Since Dudley’s integral is sharp for stationary processes (see Section 2.1 for background), (1) inherits the sharpness, even if the original process is nonstationary. The sharpness of (1) also has an interesting consequence: by essentially maximizing over $\mu$ and applying the date processing inequality, we obtain a concise new proof of the nontrivial lower bound part of the classical majorizing measure theorem. This new proof of the majorizing measure theorem is conceptually simple, as it closely mirrors the argument used for the Dudley integral upper bound, essentially reversing its key steps. It may also be interpreted as saying that Fernique’s result about sharpness of Dudley’s inequality for stationary processes [17] already contains, in essence, Talagrand’s majorizing measure theorem [37], once combined with relatively standard information-theoretic techniques. Figure 1 illustrates the implication relations of these results.
@@
Figure 1: Arrows indicate implications of results, and the techniques/assumptions for the implications. For Gaussian processes, the majorizing measure theorem and the proposed rate-distortion integral are two-sided bounds, whereas Dudley’s integral is only an upper bound unless stationarity is assumed.
@@
Second, our approach makes localization of generalization bounds more straightforward. Localization refers to the phenomenon where the index $Z$ of the process concentrates on a subset much smaller than the full index set $T$, so that taking the supremum over all of $T$ may significantly overestimate the typical value of the process at $Z$. For example, in empirical risk minimization (ERM), $Z$ denotes the estimator, which tends to concentrate near zero as the sample size increases. Traditionally, localization is achieved by the “slicing” (or “peeling”) argument, which yields sharp convergence rate for some empirical risk minimization problems [44]. Slicing typically involves controlling the modulus of continuity of the process through careful controls of the deviation magnitude for different slices. In contrast, using properties of the rate-distortion function, we can often directly upper bound the right side of (1) using the variance of $Z\sim\mu$, which is much smaller than the size of $T$. Similarly, tail probability bounds on $X_{Z}$, based on the moment generating function of $\|Z\|_{2}$, can also be derived using a lifting technique. This new method of localization does not require slicing.
To demonstrate our approach, we obtain new results for mean estimation via ERM when the parameter set is the image of the Sobolev ellipsoid under a Lipschitz map. Sobolev ellipsoid play a fundamental role in nonparametric statistics [40], whereas the Lipschitz map may represent a nonlinear feature map and is a technically challenging ingredient that necessitates the majorizing measure theorem machinery (see sub-Gaussian comparison (10)). Assuming smoothness parameter $\beta>1/2$, the traditional slicing method based on Dudley’s integral [41] yields an upper bound of order $\tfrac{1}{2\beta-1}n^{-2\beta/(2\beta+1)}$ for the mean squared error. In contrast, our method achieves the sharper rate $\tfrac{1}{\sqrt{2\beta-1}}n^{-2\beta/(2\beta+1)}$, which is beyond previous techniques (when the Lipschitz map is involved). Note that the factor $\tfrac{1}{2\beta-1}$ also appears in Dudley’s integral for the supremum of a Gaussian process indexed by the Sobolev ellipsoid, a classical example of the looseness of Dudley’s bound [39], where the optimal dependence is order $\tfrac{1}{\sqrt{2\beta-1}}$.
In [arXiv:2508.18682v2], we further present lifting techniques for tail probability bounds, and application to sparse regression, recovering results of Raskutti et al. [35] for $\ell_{q}$ balls, and further extending the rates to weak $\ell_{q}$-balls. The results for weak $\ell_{q}$-balls appears to be beyond prior techniques.
Further related work: Connections between ERM and upper bounds of stochastic processes, especially characterizing the minimax rate as the solution to a fixed point equation involving the localized Gaussian complexity, can be traced back to [41]. In the random design setting where the symmetrization argument applies, the localized Rademacher complexity was developed by [31, 4, 19]. An alternative approach for bounding the ERM risk using anti-concentration (small ball probability) rather than concentration was developed in [21, 34], which can be used in cases beyond sub-Gaussian noise. On the other hand, for convex regression classes with Gaussian noise, concentration has been used to show that the mean-squared error (MSE) is tightly bounded by the solution to an equation involving a localized Gaussian width [7, 45]. These latter methods crucially rely on convexity of the regression class, hence do not apply to some statistical examples we consider (see Remark 2).
Talagrand’s majorizing measure theorem [37] provides a sharp “geometric” characterization of the supremum of a Gaussian process in term of the $\gamma_{2}$ functional (see definition in (9)). Talagrand also developed several equivalent characterizations of $\gamma_{2}$, including admissible nets, tree-packing and labeled nets; see [39, 42] and the references therein. In particular, the labeled net (called “index” of an atom in the original paper [38, Theorem 2.1]) is closely related to the length of a uniquely decodable code, hence has an information-theoretic interpretation. This connection was clarified and further explored through the Kraft-McMillan inequality by [32, 9]. However, none of these prior works established (1). While general bounds for empirical processes and PAC-Bayes were derived based on $\gamma_{2}$ or its related forms [34, 3, 20], its application to computing minimax rates for specific classes appears to be limited, since $\gamma_{2}$ can often be more difficult to compute than directly evaluating the supremum of a process [38].
@@
A closely related line of research, started by [36] and further developed by other authors including [47, 16, 2, 5, 1], proved information-theoretic generalization bounds in terms of the mutual information between the random index and the process. A basic version (see for example [36]) is
$\mathbb{E}[X_{Z}]\lesssim\sup_{t\in T}\|X_{t}\|_{G}\sqrt{I(Y;Z)}$ (2)
@@
where the exact sub-Gaussian norm $\|\cdot\|_{G}$ will be defined in (4), and $Y:=(X_{t})_{t\in T}$ denotes the stochastic process (in the case of finite $T$, we have $Y\in\mathbb{R}^{|T|}$). While this can be interpreted as a strengthening of a maximal inequality [18], further refinement using the chaining argument, strengthening the Dudley integral, is also possible [2]. This type of results are useful in statistics and machine learning, where $Z$ represents the output of an algorithm which has limited mutual information with the training data, potentially much smaller than the covering number of $T$. Though also concerning a random index framework, their results differ from and are not directly comparable with (1), since their strengthening is in terms of the mutual information between $Z$ and the data, rather than the rate-distortion function of $Z$. Their proofs typically involve information-theoretic arguments such as chain rules and nonnegativity of the relative entropy, but not the lifting technique in the present paper. Nevertheless, we will show in Theorem 2 a result that encompasses both (1) and (2), highlighting the versatility of the lifting technique.
Finally, another direction of extending the theory about supremum of processes, considered recently by [28, 27, 10], generalizes the supremum to a soft-max, also known as free energy, which recovers the supremum as the inverse temperature tends to infinity. It seems an interesting future work to extend the lifting approach to bound the soft-max, and we reported preliminary reports in [27].
Organization: Section 2 recollects standard facts about upper bounds for stochastic processes, its applications to ERM, and relevant information-theoretic tools. Section 3 presents the main results: the rate-distortion integral, its sharpness, and an example of application to ERM over the Sobolev ellipsoid under a 1-Lipschitz map. Section 4-5 prove the upper and lower bounds for the rate-distortion integral. Section 6 presents an equivalent form of the rate-distortion integral which yields a new proof of the majorizing measure theorem, and Section 7 further applies it to the ellipsoid example.
@@ 2 Preliminaries
Notation: $\wedge$ and $\vee$ to denote the max and min of two numbers, and $x]_{+}:=x\vee 0$. Given probability measures $P$ and $Q$ on a metric space $(T,d)$, $W_{p}(P,Q):=\sup_{P_{XY}\in\Pi(P,Q)}\mathbb{E}^{\frac{1}{p}}[d(X,Y)^{p}]$ denotes the Wasserstein distance, where $\Pi(P,Q)$ denotes the set of couplings. For brevity, $\Pi(P):=\Pi(P,P)$. Product measures are denoted by $P\times Q$ or $P^{\otimes N}$. The diameter of a set $T$ is ${\rm diam}(T):=\sup_{x,y\in T}d(x,y)$. $\lesssim$ means less than or equal to up to a universal constant, and $\asymp$ means both $\lesssim$ and $\gtrsim$. We denote by $\ell_{2}$ the Hilbert space of square-summable infinite sequences, equipped by the norm $\|\cdot\|_{2}$, abbreviated as $\|\cdot\|$. The empirical distribution of a sequence $x=(x(1),x(2),\dots,x(d))$ is denoted by $\widehat{P}_{x}$. Given a set $T$ in $\mathbb{R}^{d}$ and a matrix $A\in\mathbb{R}^{n\times d}$, define $T-T:=\{t-s\colon t,s\in T\}$.
@@ 2.1 Upper bounds for stochastic processes
Classically, Dudley’s integral provides upper bounds on a sub-Gaussian process in terms of the covering numbers. Following [43, Exercise 2.40], we define the sub-gaussian variance of a random variable $X$ as
${\rm var}_{G}(X):=\inf\{\sigma^{2}\colon\mathbb{E}[e^{\lambda(X-\mathbb{E}X)}]\leq e^{\sigma^{2}\lambda^{2}/2}\textrm{ for all $\lambda\in\mathbb{R}$}\},$ (3)
and define the exact sub-Gaussian norm by
$\|X\|_{G}^{2}={\rm var}_{G}(X)+(\mathbb{E}[X])^{2}.$ (4)
This is one of the several equivalent (up to universal constant) definitions of the sub-Gaussian norm; see [43, Proposition 2.6.1]. Let $(X_{t})_{t\in T}$ be a centered (i.e., $\mathbb{E}[X_{t}]=0$) sub-Gaussian process, and define the natural metric
$d(t,t^{\prime}):=\|X_{t}-X_{t^{\prime}}\|_{G},\quad\forall t,t^{\prime}\in T.$ (5)
Then $(T,d)$ is a metric space. Dudley’ integral states that
$\mathbb{E}[\sup_{t\in T}X_{t}]\leq K\int_{0}^{\Delta}\sqrt{\ln\mathsf{N}(T,\lambda)}d\lambda$ (6)
@@
where $K$ is a universal constant, $\mathsf{N}(T,\lambda)$ denotes the $\lambda$-covering number of $T$, i.e. the minimal number of (open) balls of radius $\lambda$ required to cover $T$, and $\Delta$ denotes the diameter of $T$. While usually called Dudley’s integral, (6) is occasionally attributed to other authors as well; see [15, 33, 39] for more about the history. The Gaussian process is a special case of the sub-Gaussian process where $(X_{t_{1}},X_{t_{2}},\dots,X_{t_{n}})$ follows the Gaussian distribution for any $t_{1},\dots,t_{n}$. For Gaussian processes, a commonly referenced explicit bound is $K<24$ [22].
Recall the following definition of stationary Gaussian processes (see e.g. [42, 39]):
Definition 1.
The Gaussian process $(X_{t})_{t\in T}$ is called stationary if there exists a group $G$ acting on $T$ such that
  1. 1.
    $d(g(t),g(s))=d(t,s)$ for all $t,s\in T$, $g\in G$ (translation invariance).
  2. 2.
    For every $t,s\in T$, there exists $g\in G$ such that $t=g(s)$ (transitivity).
@@
It is well-known that Dudley’s integral is not sharp (i.e., not a two-sided inequality up to universal constants). However, for stationary Gaussian processes, it is known since Fernique [17] that Dudley’s integral is sharp:
$\int_{0}^{\Delta}\sqrt{\ln\mathsf{N}(T,\lambda)}d\lambda\leq K^{\prime}\mathbb{E}[\sup_{t\in T}X_{t}]$ (7)
@@
for a universal constant $K^{\prime}>0$. [33] provides an exposition of the stationary process story and gives a bound of $K^{\prime}\leq 432$. The stationary case is particularly convenient, since the smaller balls in the chaining argument are equivalent in each level, so that the chaining process in the proof of the Dudley integral can be reversed (see [42] for discussions).
@@
For general nonstationary Gaussian processes, the same paper of Fernique [17] conjectured what was later known as the majorizing measure theorem, whose nontrivial lower bound part was proved more than a decade later by Talagrand [37] using a more sophisticated generic chaining argument. Given a probability measure $\mu$ on $T$, define
$I_{\mu}(t):=\int_{0}^{\Delta}\sqrt{\ln\frac{1}{\mu(B(t,\lambda))}}d\lambda.$ (8)
@@
The usual definition of $\gamma_{2}(T)$ in terms of admissible partitions (developed later by Talagrand in the 90s cf. [39]) is equivalent to (9) up to a constant factor. Generally speaking, $\gamma_{2}(T)$ is more difficult to evaluate than Dudley’s integral: $\mu$ needs to be optimized for a quantity depending on $\mu(B(t,\lambda))$ over different scales of $\lambda$, whereas in Dudley’s integral, we can calculate the covering number by optimizing for each scale separately.
@@
The special case where $(X_{t})$ is also Gaussian is called the Gaussian comparison inequality, usually proved by an interpolation argument which is highly Gaussian-specific [22]. The more general sub-Gaussian case can be immediately seen from the majorizing measure theorem. As noted in [42], this version of sub-Gaussian comparison is a nontrivial consequence of the majorizing measure theorem, for which no other proof is known.
It is also known (see [39]) that the following quantities are equivalent to $\gamma_{2}(T)$ up to constants: